Thirteen specimens of 90/10 Cu-Ni alloys, each with a specific iron were tested in a corrosion-wheel setup. The wheel was rotated in salt seawater at 30 ft/s for 60 days. The corrosion was measured in weight loss in milligram/square decimeter/day, MDD. The following data were collected.

Show all work.

X(Fe) Yloss in MDD) 127.6, 124.0, 110.8, 103.9, 101.5, 130.1, 122.0,92.3, 113.1,83.7,128.0,91.4,86.2 0.01, 0.48, 0.71, 0.95, 1.19,0.01, 0.48, 1.44,0.71, 1.96, 0.01, 1.44,1.96

  1. Assuming a model, Y= β0+β1X+ε, what are the least squares estimates of β0 and β1? What is the prediction equation?
  2. Construct an analysis of variance table and test the hypothesis H0: β1=0 with an α risk of 0.05.
  3. What are the confidence limits (α=0.05) for β1?
  4. What are the confidence limits (α=0.05) for the true mean value of Y when X=3?
  5. What are the confidence limits (α=0.05) for the difference between the true mean value of Y when X1=3 and the true mean value of Y when X2=-2?
  6. Are there any indications that a better model should be tried?

(Using R is preferable for ANOVA)
Answer #1
X Y (x-x̅)² (y-ȳ)² (x-x̅)(y-ȳ)
0.01 127.6 0.7449018 352.8618 -16.2126
0.48 124 0.1545095 230.5725 -5.96872
0.71 110.8 0.0265941 3.938698 -0.32364
0.95 103.9 0.0059172 24.16101 -0.37811
1.19 101.5 0.1004402 53.51485 -2.31841
0.01 130.1 0.7449018 453.0349 -18.3703
0.48 122 0.1545095 173.8341 -5.18257
1.44 92.3 0.3214018 272.7579 -9.36295
0.71 113.1 0.0265941 18.35793 -0.69872
1.96 83.7 1.1814018 630.7825 -27.2985
0.01 128 0.7449018 368.0495 -16.5578
1.44 91.4 0.3214018 303.2956 -9.87318
1.96 86.2 1.1814018 511.4556 -24.5812
ΣX ΣY Σ(x-x̅)² Σ(y-ȳ)² Σ(x-x̅)(y-ȳ)
total sum 11.35 1414.6 5.708877 3396.617 -137.127
mean 0.873076923 108.815385 SSxx SSyy SSxy

sample size ,   n =   13          
here, x̅ =   0.873076923       ȳ =   108.8153846  
                  
SSxx =    Σ(x-x̅)² =    5.708876923          
SSxy=   Σ(x-x̅)(y-ȳ) =   -137.1266154          
                  
slope ,    ß1 = SSxy/SSxx =   -24.01989345          
                  
intercept,   ß0 = y̅-ß1* x̄ =   129.7865993          
                  
so, regression line is   Ŷ =   129.7866   +   -24.0199   *x

-----------------------------------

what are the least squares estimates of β0 and β1? What is the prediction equation?

β0 = 129.7866

and β1 = -24.0199

Ŷ =   129.7866 -24.0199   *x

-------------------------------

Construct an analysis of variance table and test the hypothesis H0: β1=0 with an α risk of 0.05

SSE=   (Sx*Sy - S²xy)/Sx =    102.85
SSR=   S²xy/Sxx =   3293.77

Anova table
variation SS df MS F-stat p-value
regression 3293.77 1 3293.77 352.2737 0.0000
error, 102.85 11 9.35
total 3396.62 12

std error ,Se =    √(SSE/(n-2)) =    3.0578
slope hypothesis test               tail=   2
Ho:   ß1=   0          
H1:   ß1╪   0          
n=   13              
alpha=   0.05              
estimated std error of slope =Se(ß1) =                s/√Sxx =    1.2798
                  
t stat =    ß1 /Se(ß1) =        -18.76895613         
                  
p-value =    0.0000              
decision :    p-value<α , reject Ho  

so, slope is not equal to zero

-----------------------------------------------------

What are the confidence limits (α=0.05) for β1

confidence interval for slope              
t critical value=   t α/2 =    2.20098516  

estimated std error of slope =Se(ß1) =                s/√Sxx =    1.2798    
margin of error ,E=       tα/2*std error =        2.8167485
lower confidence limit =        ß̂1-E =   -26.83664
upper confidence limit=       ß̂1+E =   -21.2031  

-------------------------------------------------------

What are the confidence limits (α=0.05) for the true mean value of Y when X=3?

X Value=   3
Confidence Level=   95%
  
Sample Size , n=   13
Degrees of Freedom,df=n-2 =   11
critical t Value=tα/2 =   2.201
X̅ =Σx/n = 0.873
Σ(x-x̅)² =Sxx   5.709
Standard Error of the Estimate,Se=   3.058 (calculated earlier)
h Statistic = (1/n+(X-X̅)^2/Sxx) =    0.869
Predicted Y (YHat) =
Ŷ =   129.7866 -24.0199   *3 = 57.727
  
For Average Y  
margin of error,E=t*Se*√(h stat) =    6.2751
Confidence Lower Limit=Ŷ -E =   51.4519
Confidence Upper Limit=Ŷ +E =   64.0020

-----------------------------------

only four parts can be answered per post as HOMEWORKLIB RULES answering guidelines